OpenStudy (anonymous):

what is the x and y int f(x)=x^2 − 3x − 40 / x+ 5

OpenStudy (anonymous):

Plug \(x=0\) to find the y-intercept: \(y=\frac{0-0-40}{0+5}=-8\). So, the y-intercept is \((0,-8)\).

OpenStudy (anonymous):

To find the x-intercept, plug \(f(x)=0\): \(0=\frac{x^2-3x-40}{x+5} \implies x^2-3x-40=0 \implies (x-8)(x+5)=0 \implies x=8,-5\). So, we have two x-intercepts, which are \((8,0) \text{ and } (-5,0).\)

OpenStudy (anonymous):

anwara is back!

OpenStudy (anonymous):

Hey! :D

OpenStudy (anonymous):

with some snappy latex too. good to see you again :)

OpenStudy (anonymous):

Haha, thanks! Good to see you too.

OpenStudy (anonymous):

thnx